Solve.
1/8 + 3/16x = 3/8
I came up with x=4/3 but that wasn't right.

Answer by mananth(16946)

1/8 + 3/16x = 3/8
3/16x=3/8-1/8
3/16x=2/8
24=32x
x=24/32
x=3/4